For the following cost function TC = a + bQ – cQ2 + dQ3, determine the equations for TFC, TVC, AFC, AVC, AC, MC. (a,b,c,d> 0)


1
Expert's answer
2022-03-30T17:45:34-0400

"AC=\frac{TC}Q=\frac aQ+b-cQ+dQ^2,"

"MC=TC'=b-2cQ+3dQ^2."